Here is a bivariate data set.

x y
97 45.2
33.2 95.3
58.3 59.5
58.3 84.2
61.9 41.8
50.1 73.2
47.5 97.5
24.8 70.7
9.5 115.9
52.1 68.1
52.4 84.6
18 93.6
65.8 76.5
60.4 40.4



Find the correlation coefficient and report it accurate to three decimal places.
r =

What proportion of the variation in y can be explained by the variation in the values of x? Report answer as a percentage accurate to one decimal place.
r² = %
Answer #1

ΣX = 689.3
ΣY = 1046.5
ΣX * Y = 46806.18
ΣX2 = 40374.15
ΣY2 = 84838.19

r = ( ( N * \Sigma XY) - \Sigma X \Sigma Y) / (\sqrt{ (N \Sigma (X^{2}) - (\Sigma X)^{2} )(N \Sigma (Y^{2}) - (\Sigma Y)^{2} })
r = ( ( 14 * 46806.18 ) - 689.3 * 1046.5 ) / ( \sqrt{ ( 14 * 40374.15 ) - ( 689.3 )^{2} )( 14 * 84838.19 ) - ( 1046.5 )^{2} })
r = - 0.723

Coefficient of Determination
R^2 = r^2 = 0.523


Explained variation = 0.523* 100 = 52.3%


Unexplained variation = 1 - 0.523* 100 = 47.7%